How many grams of KCIO3 are needed to make 1.5 x 10^24 particles of O2

2KCI + 3O2 = 2KCIO3
1
Expert's answer
2020-04-06T11:58:27-0400

N(O2) = 1.5*1024 particles of O2


Solution:

1)

N(O2) = n(O2) * Na;

n(O2) = N(O2) / Na;

n(O2) = (1.5*1024) / (6.02*1023) = 2.4917 moles.


The equation of a chemical reaction:

2KCIO3 = 2KCI + 3O2


2)

According to the chemical equation: n(KCIO3)/2 = n(O2)/3.

n(KClO3) = 2*n(O2)/3;

n(KClO3) = (2 * 2.4917 moles) / 3 = 1.66 moles.


3)

n(KClO3) = m(KClO3) / M(KClO3);

M(KClO3) = Ar(K) + Ar(Cl) + 3*Ar(O) = 39 + 35.5 +3*16 = 122.5 (g/mol).

m(KClO3) = n(KClO3) * M(KClO3) = (1.66 moles) * (122.5 g/mol) = 203.35 g.

m(KClO3) = 203.35 g.


Answer: 203.35 grams of KClO3.